Problem 1-2 Suppose you bought 200 shares of stock at an initial price of $38 per share. The stock paid a dividend of $.30 per share during the following year, and the share price at the end of the year was $41. a. What is the capital gains yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Omit the%sign in your response.) Capital gains yield b. What is the dividend yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Omit the % sign in your response.) Dividend yield c. What is the total rate of return on the investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Omit the%sign in your response.) Total rate of return
Answer #1

a.Capital gains yield=(End price-Beginning price)/Beginning price

=(41-38)/38=7.89(Approx).

b.Dividend yield=Dividend/Beginning price

=(0.3/38)=0.79(Approx).

c.Total rate=Capital gains yield+Dividend yield

=(7.89+0.79)=8.68(Approx).